Sol Summary – April 2nd

Sol: 2
Summary Title: Chilling (or troubleshooting) on Mars
Author’s name: Alba Sánchez Montalvo
Mission Status: Ongoing
Sol Activity Summary:
The day started with a slow morning today. The team decided not to do any EVAs during this morning to rest and recover from the sleepless and frenetic days they had before arriving at the MDRS. Maxime made pancakes for breakfast and after that, some people worked on their experiments, played card games or had some time to do sports. After lunch, Louis, Loriane, Maxime and Hippolyte went on an EVA to set Maxime’s weather station and to start mapping our surroundings for Louis’ experiment. Arnaud, Romain, Imane and Alba stayed at the station and worked on their experiments. For Imane’s experiment’s, she weighed the stool tubes collected yesterday. She also weighed the next collection tubes. Hippolyte collected the first tomatoes from the GreenHab, they taste really good ! When the crew members came back from their EVA, everyone sat together to work on the reports about daily life and operations at the station before preparing dinner and having some time for team building to finish the day.
Look Ahead Plan: 2 EVAs (one in the morning and one in the afternoon) will be performed. The first one will be performed by Maxime and Alba. They will pursue Maxime’s experiment and explore a new area. The second EVA will focus on continuing Louis’ experiment and automatically flight the drone above more diverse landscapes. To help Louis in his endeavor, Hippolyte and Arnaud will be joining him in his journey. The crew members staying in the station will continue to work on their respective experiments but they will be no need to take medical samples from the crew members.
Anomalies in work: Louis (Engineer) had struggles with setting up his experiment. At first, the drone wouldn’t take off because of magnetic interference even though the drone was isolated on the ground. After a few minutes of troubleshooting with Hippolyte’s help, the drone was able to take off. During the same EVA, Maxime had problems with his weather station. His wind indicator partially broke and a future EVA should be planned to repair it.
Weather: blue and clear skies, sunny and warm the whole day!
Crew Physical Status: everyone is feeling good after some rest
EVA: They began by exiting their station to secure Maxime’s weather station to the Curiosity rover, heading towards Pooh’s Corner. However, the weather station suffered damage on the way, specifically losing its wind direction indicator. Despite this, they managed to set up the station west of Pooh’s Corner and activated its cameras. They noted a low battery alert but proceeded with their mission, planning to address these issues in a follow-up EVA.
Simultaneously, another team struggled with magnetic interference affecting their drone’s takeoff. With some troubleshooting, they overcame this obstacle, enabling the drone to complete its first automatic flight, which validated the basic parameters of their experiment.
Following their respective tasks, they gathered and headed towards Cowboy Corner for further exploration. However, due to navigational errors, they stopped near North Ridge instead, where they decided to climb the hills to gather additional data through pictures and videos. The day concluded with the team returning to their base, marking the end of their EVA.
Reports to be filed: Sol summary, Journalist’s report, Greenhab report, operations report, EVA report, EVA request, astronomy report.
Support Requested: the biomedical team made a request for Sergii to pick up a package at the end of the week and asked for permission to leave the station for a few hours during the last day of simulation to arrange the shipment of the different samples collected during the experience to Belgium.

Copyright © The Mars Society. All rights reserved. | Main Site